### v2

# ColorSense: A Study on Color Vision in Machine Visual Recognition

Abstract
Color vision is essential for human visual perception, but its impact on machine perception is still underexplored. There has been an intensified demand for understanding its role in machine perception for safety-critical tasks such as assistive driving and surgery but lacking suitable datasets. To fill this gap, we curate multipurpose datasets ColorSense, by collecting 110,000 non-trivial human annotations of foreground and background color labels from popular visual recognition benchmarks. To investigate the impact of color vision on machine perception, we assign each image a color discrimination level based on its dominant foreground and background colors and use it to study the impact of color vision on machine perception. We validate the use of our datasets by demonstrating that the level of color discrimination has a dominating effect on the performance of mainstream machine perception models. Specifically, we examine the perception ability of machine vision by considering key factors such as model architecture, training objective, model size, training data, and task complexity. Furthermore, to investigate how color and environmental factors affect the robustness of visual recognition in machine perception, we integrate our ColorSense datasets with image corruptions and perform a more comprehensive visual perception evaluation. We jointly analyze the impact of color vision and image corruption on machine perception. Our findings suggest that object recognition tasks such as classification and localization are susceptible to color vision bias, especially for high-stakes cases such as vehicle classes, and advanced mitigation techniques such as data augmentation and so on only give marginal improvement. Our analyses highlight the need for new approaches toward the performance evaluation of machine perception models in real-world applications. Lastly, we present various potential applications of ColorSense such as studying spurious correlations.
I Introduction
The development of deep neural networks (DNN) has been deeply rooted in inspiration drawn from neuroscience and cognitive science [53]. For instance, convolutional neural networks (CNN) [37, 24] have enjoyed great success in various tasks in computer vision by modeling the human visual cortex system. More recently, Transformer [61] has become a widely adopted architecture that models the human attention mechanism and allows better contextual understanding. It has since expanded its success beyond natural language processing to computer vision [17]. Some recent studies explore the behavioral aspects of DNN models for object recognition [60, 21, 29]. However, there are other important parts of the visual pathway of primates that remain underexplored, such as color vision [52, 34], a neuroscience topic that studies how color affects human visual perception. With such a nomination, we curate new datasets to study machine visual recognition as a proxy task to connect the color vision aspects of human visual perception with those of machine vision.
Figure 1: Mapping of our ColorSense datasets to human visual aspects and examples of our ColorSense integrated with corrupted images (ImageNet and CIFAR10). (a) ColorSense can evaluate the color vision and more complete visual perception capabilities. (b-c) Examples of ColorSense integrated with corrupted datasets. Each row indicates a color discrimination level.
Color is a fundamental perception that we often take for granted as humans. It is only when our color vision is hindered, as in the cases of colorblindness, that we realize its true significance. Similarly, color plays a role in machine vision [46, 20, 44] and has a crucial impact on downstream safety-critical visual recognition tasks such as self-driving, detection of skin disease, or robotic-assisted surgery. Previous research has identified similarities between human and machine vision. For example, the last convolution layers in classical machine vision models are shown to have high color sensitivity, similar to the most anterior occipital visual areas in primates [19], and strong white balance can negatively impact DNNs models in image segmentation and classification tasks [5]. However, comprehensive evaluations of cutting-edge models remain limited due to the lack of well-annotated datasets that can be seamlessly integrated with modern computer vision research. To the best of our knowledge, [14] was the only proposed dataset to date that aims to facilitate the studies on the effect of color distortion. To address this need and to enable a wider range of studies on color vision, we curate ColorSense-CIFAR and ColorSence-ImageNet. We manually separated ImageNet into eight groups (red, green, blue, gray, brown, black, white, others) based on the dominant foreground and background colors of each image. Furthermore, to study how neighboring colors help machines discriminate objects, we consult color professionals and the established standard [1] to categorize the foreground/background color pairs into 3 color discrimination groups {Hard, Medium, Easy} (details in §III). We call this dataset ColorSense, designed to quantify the sense of color for machine vision. To achieve this goal, we evaluate models beyond the regime of using total accuracy as the sole objective and examine model performance in each stratified color discrimination groups. Similarly, on CIFAR, we also provide dominant foreground color attributes (Fig. 2(b)) to create ColorSense-CIFAR. We mainly focus on ColorSense-ImageNet in this work.
As a benefit, ColorSense can also be seamlessly integrated with out-of-distribution (OOD) datasets like ImageNet-C [27] (dubbed ColorSense-ImageNet-C) as a complement to broaden the analysis of color vision to include simulated environmental factors such as noisy data shifts. For human vision, it is well-established that the primate visual system can effortlessly and robustly discriminate between various visual objects, regardless of the wide range of images that each object can produce during natural vision [16, 71, 51, 32, 55, 42, 59, 18, 54], and the robustness should extend to different surroundings or added noise [43, 52, 34]. Fig. 1 illustrates where ColorSense fits in bridging the evaluation of visual perception between humans and machines.
To gain a comprehensive understanding of color vision in machine vision, we conduct comprehensive analyses on ColorSense in conjunction with corrupted datasets across different model sizes and architectures, training objectives, etc., which shows the diverse capabilities of ColorSense. This framework provides new insights into model behaviors under different environmental and color discrimination conditions. Our study is a step toward understanding color vision in deep learning, and we believe ColorSense will facilitate benchmarking and evaluation efforts in different ways. In addition, we explore advanced bias mitigation and provide reasoning behind the phenomenon. Lastly, we showcase other uses of our dataset: we perform case studies on a safety-critical task to demonstrate the importance of color vision, quantify model robustness, and study fairness using color subgroups from ColorSense.
To summarize, our contributions are threefold:
  • We curate ColorSense dataset with 110,000 human annotations to study how color vision affects computer vision models on various visual recognition tasks.
  • We conduct comprehensive analyses on factors like model size, architecture, training objectives and procedures and propose generic metrics for the color vision effect and model robustness.
  • We show ColorSense’s diverse uses, such as integrating with ImageNet-C to study the robustness of machine color vision in natural and simulated noisy environments, extended use in additional recognition tasks like object localization, and studying spurious correlations.
II Bridging Human Vision and Machine Vision
II-A Human Vision Test
We, as humans, rely on stimuli from light that reflect on objects to our eyes to perceive objects. More importantly, our retinal cone cells respond differently to wavelength in a diverse spectrum of light [52, 34]. For instance, three different types of cone cells respond predominantly to red, green, and blue. This color vision mechanism helps us discriminate objects under different contrast, environments, etc., and is an essential part of why our vision can adapt quickly [52, 34, 25] and thus become more robust [16, 71, 51, 32, 55, 42, 59, 18, 54]. More concretely, the opponent-process theory, a theory that different wavelengths would oppress one type of cell, explains how psychological perception would affect our way of discriminating objects [52, 34]. Such phenomena have important real-life applications; for example, a red apple would be more clearly identified when placed on a green background (Fig. 1) than on an orange background; army soldiers would wear green and brown uniforms instead of white for better camouflage. To emphasize the distinction between visual responses to different colors [4], we mainly use color discrimination as our terminology in experiments.
To measure human visual perception ability, a common routine is to perform the Snellen’s test, where subjects need to discriminate black alphabets in a white background from a certain distance. This test measures our visual acuity, the ability to distinguish shapes and the details of objects [50, 47]. However, a 20/20 vision does not indicate perfect vision [50, 47, 33], and ophthalmologists have proposed the contrast sensitivity test and some color vision tests. The former measures the ability to discriminate finer and finer increments of light versus dark [26], and the latter includes color plate test, hue test, anomaloscope test [2] to measure the accuracy of color vision.
II-B Machine Vision Test
Unlike various early visual perception tests for human vision, machine vision often goes directly to recognition tasks. Although DNN models reach human-level performances on diverse visual recognition tasks [64, 63, 65], they are not always robust [23, 27, 10], which means there are notable gaps to true human-aligned vision. Hence, instead of revolving around performances on generalization errors, we take a step back and ask: What are the fundamental aspects that current studies overlooked? Inspired by the neuroscientific literature, our answer is to study how color vision affects machine visual recognition, as an important first step to bridge the gap. Prior works have touched on the related contrast sensitivity concept [38, 6] and Olah et al. [46] visualizes early CNN layers to hint color’s role. But due to the lack of proper datasets, they did not evaluate DNNs’ color vision behavior that affects visual recognition.
III Benchmarking for Color Vision
In the section, we introduce our ColorSense-ImageNet, a dataset that are based on the validation sets of ImageNet [15]. On ImageNet validation set, we label 50,000 images’ dominant colors in both the foreground and background. To enable color vision study in different setups, we also provide 10,000 color labels on CIFAR10 [36], forming ColorSense-CIFAR10. With our labeling efforts, we can create color discrimination groups for measuring color discrimination ability in computer vision and visual recognition. In the following sections, we mainly focus on the evaluations of ColorSense-ImageNet and leave CIFAR10 results in the Appendix.
III-A Labeling Process
  1. 1.
    We label the perceived color that has the most coverage on an object. For example, in Fig. 2 (c-top), the car is almost red everywhere so we label foreground as “red.”
  2. 2.
    When two or three colors have significant coverage, we choose the color that has more coverage. We label the example in Fig. 2 (c-mid) as “brown”, though the black track is also significant.
  3. 3.
    When the perceived color does not belong to our categories, or when multiple colors appear on the object and none is larger than the rest, we put it in the “others” category (Fig. 2 (c-bottom)).
A similar procedure is applied to label the dominant background color.
Figure 2: Examples of our ColorSense datasets and labeling principles. (a) Top two rows are ColorSense-ImageNet-Foreground, and the bottom two rows are ColorSense-ImageNet-Background. In conjunction, we can define color discrimination groups (§ III). (b) ColorSense-CIFAR10. (c) Examples of the three labeling principles. See § III-A for full labeling details.
III-B Defining Color Discrimination Groups
We follow the Web Content Accessibility Guidelines (WCAG) [1] and consult experts from the School of Fine Arts to assign all possible color pairs into three color discrimination (CD) groups, {Hard, Medium, Esay}. The Hard groups consist of images with the same foreground and background colors. The Easy group includes color pairs with a score larger than the WCAG Minimum Contrast Success Criterion 1.4.3 [1] and they also correspond to the top one-third scores ranked among all pairs; the rest goes into Medium. Tab. V (in Appendix) presents the statistics for the number of color pairs in each CD group and the number of test images in each group for ColorSense datasets. We discuss the label quality and provide details of the color pairs for each color discrimination group in the Appendix.
IV Evaluation: Image Classification
IV-A Fundamental Questions
Does color vision affect DNN models as it affects humans?
While DNNs are artificial models of neurons in the brain, some of them are designed to mimic the structure and function of the visual cortex [37]. As such, it is possible that these models possess some form of color discrimination, similar to how humans perceive color. This hypothesis raises the question of whether color vision affects DNN models and how this could impact their performance in various applications. Further research is needed to explore the relationship between color vision and DNN models and to uncover any potential benefits or limitations. Prior works show most humans are robust to color vision groups, but the decisions vary in response time (we conduct a small-scale human study to re-verify it in the Appendix) [16, 71, 51, 32, 55, 42, 59, 18, 54, 21], whereas for machines, the distinction could be reflected in the form of performance difference.
Does architecture matter?
The choice of deep learning architecture can have a significant impact on the performance and efficiency of a model [41, 24, 39, 40, 57]. CNNs, such as ResNets, have been the go-to architecture for computer vision tasks, thanks to their ability to learn spatial features from images. However, the emergence of newer architectures, such as the ViTs, has challenged this dominance. Unlike traditional CNNs, ViT uses self-attention mechanisms to capture global image features, allowing it to outperform CNNs on some tasks and claim less inductive bias. The debate over whether architecture matters and which one to choose, continues to be an active area of research and development. We compare diverse architectures to answer this question and hope to shed light on this area in terms of color vision.
Does model size matter?
The relationship between model size and robustness is an intriguing topic of research in the field of deep learning. Hendrycks et al. [28] showed that larger models tend to be more robust to added noises. This finding is particularly interesting since larger DNNs do resemble the number of neurons in the human brain more closely. Furthermore, human vision is known to be robust to color variations, which raises the question of whether larger DNNs are also more robust to color variations. Exploring the relationship between model size, robustness, and color variation could lead to useful insights into the functioning of DNNs and their potential applications. We compare (1) architectures of the same family and (2) models of similar sizes but different architectures to answer this question in our problem setup.
Figure 3: Testing models with different architectures and model sizes (ImageNet). We observe similar trends in the three CD groups (Hard, Medium, Easy) across all models, showing the effect of color vision is universal across architectures and model sizes.
IV-B Main Findings
To answer the three fundamental questions in Sec. IV-A, we conduct experiments with a diverse set of models trained on ImageNet (model weights are fixed from PyTorch). To account for the wide range of application setups, we select both classic and modern CNNs such as Resnet [24], Resnext [68], Convnext [41] and MobileNetV2 [57], and vision transformers such as ViT [17] and Swin Transformers [39, 40].
DNNs are deeply affected by color vision
We observe a consistent performance increase from the Hard CD group to the Easy group (Fig. 3) across all the aforementioned architectures. We perform Paired t-Tests between the CD groups across all models. We pool the Hard group accuracies from all models and Easy group accuracies from all models and then performed the statistical tests on these two sets of data. The p-value for Hard-Medium is $1.39e$-$6$, Hard-Easy is $8.34e$-$7$, and Medium-Easy is $9.87e$-$6$, which are all statistically significant. We also perform a Page’s Trend Test for increasing means ($H_{0}:\mu_{Hard}=\mu_{Medium}=\mu_{easy},H_{A}:\mu_{Hard}<\mu_{Medium}<\mu_{Easy}$) and the p-value is $5.95e$-$7$, which is also statistically significant. These results suggest that the DNNs do have color vision ability-related issues, and color discrimination abilities should be taken into account in the future when designing and evaluating computer vision models. Second, it suggests that there is room to improve the performance of DNNs on color-based tasks by developing models that can better discriminate between colors. These implications point out a deficiency of DNNs models and shed light on an area of focus for the improvement of next-generation computer vision applications.
Model size and architecture do not add obvious robustness to color vision
For each model architecture, we evaluate its multiple variations based on size. To our surprise, Fig. 3 shows that the performance gaps between the Easy and Hard groups do not obviously decrease as model size increases, suggesting the impact of the color discrimination group remains almost unchanged. In addition, we compare model architectures while controlling for model size; for example, models with 88M parameters (Convnext_base, Resnext101, SwinV2_b, ViT16_b) show similar performance gaps. While model size and architecture are essential factors to consider when designing and evaluating DNNs for various computer vision tasks, they may matter less when it comes to color vision discrimination — we observe an upward trend in performance towards the Easy group regardless of the architecture or size of the model. These suggest that human-aligned color discrimination ability may not be inherent in DNNs, and developing specialized models or techniques to address color vision discrimination may be necessary.
IV-C Advanced Analyses and Ablations
We have identified the effect of color vision in DNNs. To comprehensively understand its relationship with DNNs and potential mitigation, we consider additional factors across the entire deep learning pipeline, encompassing advanced training methods and training data, and evaluate foundation models like CLIPs. Further, we explore the color vision effect in real-world scenarios such as high-stake applications and environmental factors. We also propose metrics to evaluate model robustness to color vision effect. Finally, to identify potential causes of the phenomenon, we perform controlled experiments on luma and chroma, and conduct frequency analysis.
Advanced training setups improve color vision bias marginally
Recent works have shown that specific training recipes, such as a mix of long training epochs and complex data augmentations, can enhance model robustness and mitigate model biases [11, 62, 67, 12, 70, 69]. However, our results suggest no significant changes in overall robustness to color discrimination groups (Fig. 4), as measured by absolute gap (AG), $AG=|acc_{max}-acc_{min}|$, where $acc_{max/min}$ is the maximum/minimum accuracy of the three groups (Tab. I). Advanced data augmentation and training only reduce AG by an average of 0.58% in accuracy. This bolsters our belief that the color vision phenomenon exists in machine vision.
Figure 4: Color vision performance of standard and advanced procedure-trained models. Advanced training only makes models slightly more robust to color discrimination (also see Tab. I).
TABLE I: Average absolute gap in accuracy (%) by architecture type and dataset.
Absolute Gap MobileNet MobileNet_adv Resnet Resnet_adv Resnext Resnext_adv ColorSense 2.86 2.62 3.12 2.21 2.97 2.56 Land vehicle-only 4.35 4.87 6.59 4.27 4.63 5.87 ColorSense-C 3.37 n/a 4.40 n/a 4.46 n/a Absolute Gap Convnext SwinV2 ViT ColorSense 2.17 2.18 2.48 Land vehicle-only 3.40 4.31 6.93 ColorSense-C 4.59 4.40 4.64
Changing training data does not alter the color discrimination phenomenon
To see if color discrimination can be observed when using alternative training data, we include the results of ViTs pretrained with ImageNet21k. We observe that on ViTs, the color vision effect is present (Fig. 5-left).
Figure 5: Color vision in effect no matter how models are pre-trained. Left: ViT accuracies (%) pre-trained on ImageNet-21k. Right: CLIP 0-shot performances (%) with different training data size.
Consistent patterns from self-supervised models
We evaluate on self-supervised models like the DINO with four variants different in sizes (ViT-S/16, ViT-S/8, ViT-B/16, ViT-B/8, in this order) and report their accuracies on ImageNet in Tab. II. We observe a consistent pattern as in previous sections in these self-supervised models.
TABLE II: DINO evaluations on ImageNet
CD Group ViT-S/16 ViT-S/8 ViT-B/16 ViT-B/8
Easy 77.90 80.58 78.78 80.90
Medium 76.45 79.25 77.71 79.41
Hard 75.11 77.79 76.40 78.60
Zero-shot Foundation Models behave similarly again
We include CLIP zero-shot classification results (Fig. 5-right) for ViT-B32 and ViTB14 with two different training data: Laion400M and Laion2B. We observe a general trend: color vision is still in effect. Together with advanced training and changing training data results, the three analyses suggest that training procedures do not alter the phenomenon.
High-stakes examples have larger gaps
Color vision is important for humans to perform high-stakes tasks such as driving or cardiovascular surgery. Prior work has shown potential lighting changes can lead to dangerous erroneous behavior [49]. With the popularity of self-driving cars and the previously established knowledge about color vision in machine vision, we apply our analysis specifically to the vehicle classes in ImageNet. Ideally, we want models to be the least affected by color vision on the car-related subset. To our surprise, we observe models generally have even larger gaps (p-val = 4.3e-2) with average AG difference = 2.72% (Tab. I & Fig. 7) between CD groups. It suggests vehicle classes are affected by color vision even more, and surprisingly, some larger models, such as ViT_L is affected more than ViT_b. Advanced training (Tab. I) only reduces the color vision gap for Resnets but increases for Resnexts and MobileNetV2. Our evaluations call for further color vision studies for high-stakes tasks.
Figure 6: Color discrimination capabilities with added noises by models and by Gaussian noise (bottom-right). Models of the same total performances can have very different color vision behaviors in the presence of different corruptions. Bottom-right: CNNs see less obvious gaps between Medium and Easy group than transformers when presented Gaussian noise. These phenomena show that each architecture is built differently and is suitable for different scenarios.
Environmental factors play a role
Real-world applications of DNNs often also involve environmental factors. A benefit of ColorSense is the easy incorporation with OOD datasets like ImageNet-C to task models with added environmental factors and noises. We observe a consistent trend: color vision plays an important role in all types of corruption. In Fig. 6, we show three models, Convnext_s, Swin_b, SwinV2_s of similar accuracies on ImageNet, 83.61, 83.58, and 83.71, respectively (less than 0.15 difference) for comparisons. Even though the generalization performances are almost the same, for each corruption type, the color vision performance is different. For instance, Swin & SwinV2’s color discrimination capabilities are better than Convnext_s for all corruptions but jpeg compression, and SwinV2_s is slightly better than Swin_b on digital corruption such as pixelate but worse on noise corruptions. These findings have implications and insights for decisions such as which model we should choose? If we expect noisy environments for some custom application, we should choose SwinV2_s model among the three.
Also, we conduct case studies on specific corruptions, such as Gaussian noise (Fig. 6-bottom-right) due to its commonality in image processing. We observe that CNNs generally have less obvious gaps in Medium and Easy than ViTs. This analysis reveals the unique advantage of CNNs over ViTs. Similar analyses can be done for other corruptions. One can easily use our analysis framework to assess different models in more realistic settings, considering the effect of color vision.
Figure 7: Comparisons of all vs. land vehicle classes on ImageNet. Overall larger gaps between Hard and Easy group on vehicle classes suggest potential safety concerns when deploying such models on the road.
Figure 8: Proposed sCR & sCVR metrics. Overall quantified color vision robustness to color vision effect. Symbol sizes are proportional to model sizes.
Quantifying overall model robustness to color vision effect
To jointly consider the color vision effect and performance on corrupted dataset, we propose two metrics: scaled color vision robustness (sCVR($\uparrow$)), $\frac{acc_{tot}}{\sigma_{w}}$, where $acc_{tot}$ is the total accuracy on ImageNet-C and $\sigma_{w}=\sqrt{\frac{\sum_{i=Easy}^{Hard}w_{i}(acc_{i}-\bar{acc})^{2}}{\sum_{i=Easy}^{Hard}w_{i}}},$ is the weighted standard deviation across CD groups. Similarly, we define scaled corruption robustness (sCR($\uparrow$)), $\frac{acc_{tot}}{\sigma_{s}}$, where $\sigma_{s}=\sqrt{\frac{\sum_{s=1}^{5}(acc_{s}-\bar{acc})^{2}}{\sum_{s=1}^{5}w_{s}}}$ (across five severity levels). Fig. 8 shows the sCVR and sCR of each model. Per our metrics, Convnext_l is the overall best. We also tried the $\frac{acc_{tot}}{AG}$ as a metric, and the results were similar. We observe that most models have better sCVR as their size grows. However, for ViTs, as also observed in the land vehicle classes, there is an inverse correlation to model size, which is another anomalous model behavior our dataset reveals. We discuss the practicality of our metrics in the Appendix.
Next, we explore the potential causes of the color vision effect.
Less color vision bias on grayscaled images
To test if chroma is the main cause of the color vision effect, we perform analysis on grayscaled ImageNet. We quantified the absolute gap between the Hard and Easy groups for color and grayscale images, Fig.9-left, and observe larger gaps (more pronounced color vision effect) in color images (p-value = 3.9e-3), suggesting a role of chroma in color vision effect. However, we still observe lower performance in harder CD groups (Fig.9-right), suggesting the existence of other non-trivial causes.
Bias persists after controlling luma
To test if lumination condition lead to the observed color vision effect, we subset images with average luma in certain ranges like $(-0.2,0.2)$ $(-0.4,0.4)$ and then conduct the same analyses. Table III shows that when luma is similar, we still observe similar color vision trend (p-val = 5.95e-7 for range (-0.2, 0.2); p-val = 5.95e-7 for range (-0.4, 0.4)), suggesting that chromatic differences play a major role when lumination is controlled.
Figure 9: Left: Absolute gap (AG) comparisons. The discrimination effect is more pronounced in color images, supporting our primary finding. Right: Grayscale ImageNet results (%). We still observe slight color vision effects between the CD groups.
TABLE III: The color vision effect remains after controlling luma. Results reported in Hard/Medium/Easy CD group format.
Luma range (-0.2,0.2) (-0.4,0.4) Resnet50 73.59/77.05/77.91 73.77/76.51/77.47 Resnet101 74.87/77.60/79.75 75.20/77.46/79.09 ConvnextT 80.83/82.66/83.17 81.14/82.54/83.21 ConvnextB 82.32/84.20/85.21 82.47/83.75/85.06 VitB16 79.30/80.89/82.34 79.33/80.85/82.20 VitL16 78.91/79.65/80.85 78.37/79.83/80.93 SwinV2T 80.09/82.16/82.95 80.08/82.19/82.85 SwinV2B 82.22/84.31/85.11 82.65/84.07/84.97
TABLE IV: Object localization accuracy. The color vision effect exists in another visual recognition task like object localization, not only in classification.
CD Group IoU-30 IoU-50 IoU-70 Easy 75.31 73.12 50.34 Medium 74.48 72.77 50.01 Hard 72.94 70.80 48.97
Intuition on root cause from frequency analysis
Inspired by [30], we hypothesize that color vision bias arise from certain frequency components of the images. We conduct frequency analysis and split the frequency components with $r=$ [8, 16, 32]. We split each image into high and low frequency components (HFC/LFC), and test Resnet50 and ViT_L on the two components respectively. Results for each CD group are summarized in Tab. LABEL:tab:frequency. We observe that when presented with only the LFC of images, models consistently perform worse in Hard group, and surprisingly, on HFC both models show better performance for Hard group. Since LFC results resemble what we observe for the original images, we hypothesize that the color vision bias is predominately driven by the LFC in the images. Furthermore, as we already observed lower performance in harder groups on the grayscale experiments (Fig. 9) and considering LFC usually contains most of the color information [22], we conjecture that the color vision bias may be relevant to other low-level features that models pick up, beyond mere colors. Consequently, mitigating this bias is non-trivial, as also supported by experiments from previous sections (Tab. I & Fig. 4 & Fig. 5), which also substantiate the merit of our dataset as this phenomenon cannot be explored before.
Figure 10: Left: Object localization performance on ImageNet (threshold = IoU-30) — vehicle classes are more affected by the color vision effect. Right: Background and foreground colors as spurious correlations; green objects/background are favored by Resnet50.
Qualitative results on failure cases
We present nine common failure cases across multiple state-of-the-art models such as ConvNeXt and Swin-Transformer from the HARD group in Fig. 11. These failure cases are deliberately selected to represent scenarios where human visual recognition remains robust, yet the models exhibit significant shortcomings.
Accompanying each image is the corresponding attention map, we observe that the model may not be able to attend to the right object areas in whole, or sometimes the object is blended in the background. Thus, we believe that the models’ misclassifications stem primarily from the small color difference between foreground and background elements, which appear to impede the models’ ability to effectively discern and isolate salient features. This phenomenon underscores the ongoing challenges in developing robust computer vision systems capable of matching human perceptual acuity across diverse visual scenarios.
Figure 11: Examples misclassified by models their respective attention maps. Though the objects are easy for humans to recognize, we believe the models’ failure stem primarily from the small difference in background and foreground colors.
V Other Utilities: Object Localization and Spurious Correlation
Our ColorSense has enormous potential for a wide range of applications in computer vision. We already show its usefulness for perception task such as image classification, where color plays a critical role in recognizing objects in an image. We conduct another key visual recognition task, object localization (OL). Other than that, we can study color as a spurious correlation to analyze subgroup performance. In general, ColorSense is a valuable tool for any task that studies color discrimination in computer vision.
Object localization is similar to classification
We apply the same analysis procedure in Sec. IV-B to object localization [8]. We report top-1 localization accuracy with three different IoU thresholds (30, 50, 70). We summarize the results on ColorSenese-ImageNet in Tab. IV. We observe similar performance differences as in classification task, where the model tends to perform better as the CD group becomes easier. And this trend happens across all three different IoU thresholds. This again bolsters our finding of the color vision effect in machine vision.
High-stakes examples in object localization
Similar to classification, we conduct a case study on land vehicle classes for OL due to their resemblance to a self-driving setup. We also observe larger AG between CD groups, meaning an amplified effect of color vision. Fig.  10-left shows the IoU-30 result and we observe similar trends using other thresholds. We emphasize the worst-case situation in high-stakes tasks. For the Hard group, we observe worse performances on vehicle classes than all classes in all three thresholds, suggesting higher safety risks in those conditions and the necessity of risk mitigation. We believe our findings can be extended to other self-driving studies.
Background and foreground color as spurious correlations
Another application of our dataset is to leverage the background/foreground labels as spurious contextual correlations. For example, on ResNet50, we observe “subgroup degradation,” uneven performances across subgroups, for both foreground and background. For instance, ResNet50 tends to bias in favor of green objects and objects in a green background (Fig. 10-right). Also, we observe similar trends across all models, showing the universality of this bias. The complete plot for all models is shown in Appendix.
VI Discussion
Our work presents the development and demonstrates the utilities of our curated ColorSense dataset, which is a versatile tool for (1) investigating the impact of color vision on various visual recognition tasks, (2) featuring compatibility with OOD environmental and artificial noise datasets, and (3) controlling for spurious correlations. We argue that current literature on machine vision evaluations does not fully explore the role of color vision in object recognition and show that our framework provides a more comprehensive approach to evaluating machine visual recognition.
Our work demonstrates that current machine vision is susceptible to color vision bias, regardless of DNN architecture, training methods (e.g. self-supervised, contrastive, source data, etc.), size of the model, or controlling factors like luma. We provide quantitative evidence that machine vision is affected by color vision in the form of performance reduction in harder CD groups, and also show that advanced training procedures for robustness enhancement only marginally mitigate the color vision effect. In addition, qualitative examples suggest the background and foreground similarity impedes correct classification when the examples are easy for humans. Moreover, we reason and provide intuition about such a non-trivial phenomenon with frequency analysis, which suggests that more innovations are needed to improve model robustness to color to align with human vision. Further, in the high-stakes use cases studied, the color vision effect may be more obvious. Lastly, CNNs such as Convnext can be as powerful as transformers, as similarly observed in [58].
We believe that in the era of foundation models with various scaling laws, our dataset and analysis framework are valuable and can push the boundaries of benchmarking.
VII Potential Broader Impact
The exploration of machine-human alignment stands as a highly coveted focus in research [3]. It is anticipated that advancements in our understanding of human vision [43, 13, 56] will pave the way for notable breakthroughs in deep learning, specifically addressing the color vision bias in models. Our ColorSense dataset is positioned as a versatile and valuable resource for the comprehensive assessment of newly proposed models and training methods [7, 9]. As demonstrated in previous sections, this dataset can not only be used to study model biases but also contribute to heightened awareness of safety, spurious signals, and fairness concerns, thereby enriching the field of Machine Learning community.

### v3

@@
 Lastly, we present various potential applications of ColorSense such as studying spurious correlations.
+[^1] Dataset is available at https://github.com/charismaticchiu/ColorSense [^2] This work has been accepted for publication in the IEEE Conference on Secure and Trustworthy Machine Learning (SaTML).
+The final version will be available on IEEE Xplore.
 I Introduction
@@
 Tab.
-V (in Appendix) presents the statistics for the number of color pairs in each CD group and the number of test images in each group for ColorSense datasets.
+VI (in Appendix) presents the statistics for the number of color pairs in each CD group and the number of test images in each group for ColorSense datasets.
 We discuss the label quality and provide details of the color pairs for each color discrimination group in the Appendix.
@@
 [28] showed that larger models tend to be more robust to added noises.
-This finding is particularly interesting since larger DNNs do resemble the number of neurons in the human brain more closely.
-Furthermore, human vision is known to be robust to color variations, which raises the question of whether larger DNNs are also more robust to color variations.
+Furthermore, human vision is known to be robust to color variations [16, 71, 51, 32, 55, 42, 59, 18, 54, 21], which raises the question of whether larger DNNs are also more robust to color variations.
 Exploring the relationship between model size, robustness, and color variation could lead to useful insights into the functioning of DNNs and their potential applications.
@@
 In Fig.
-6, we show three models, Convnext_s, Swin_b, SwinV2_s of similar accuracies on ImageNet, 83.61, 83.58, and 83.71, respectively (less than 0.15 difference) for comparisons.
+6, we show three models, Convnext_s, Swin_b, SwinV2_s of similar accuracies (%) on ImageNet, 83.61, 83.58, and 83.71, respectively (less than 0.15 difference) for comparisons.
 Even though the generalization performances are almost the same, for each corruption type, the color vision performance is different.
@@
 Luma range (-0.2,0.2) (-0.4,0.4) Resnet50 73.59/77.05/77.91 73.77/76.51/77.47 Resnet101 74.87/77.60/79.75 75.20/77.46/79.09 ConvnextT 80.83/82.66/83.17 81.14/82.54/83.21 ConvnextB 82.32/84.20/85.21 82.47/83.75/85.06 VitB16 79.30/80.89/82.34 79.33/80.85/82.20 VitL16 78.91/79.65/80.85 78.37/79.83/80.93 SwinV2T 80.09/82.16/82.95 80.08/82.19/82.85 SwinV2B 82.22/84.31/85.11 82.65/84.07/84.97
-TABLE IV: Object localization accuracy. The color vision effect exists in another visual recognition task like object localization, not only in classification.
+TABLE IV: High & low frequency components accuracy (%). Results reported in Resnet50 / ViT_L format.
+LFC Easy Medium Hard $r=8$ 4.04 / 9.64 3.44 / 9.12 3.23 / 8.04 $r=16$ 21.47 / 36.49 19.27 / 34.60 17.81 / 31.39 $r=32$ 51.71 / 64.43 49.01 / 62.52 47.48 / 60.61 HFC Easy Medium Hard $r=8$ 44.97 / 51.08 44.62 / 51.61 45.52 / 52.19 $r=16$ 13.99 / 9.58 14.65 / 10.20 15.00 / 11.23 $r=32$ 2.01 / 1.82 1.85 / 1.64 2.57 / 2.67
+TABLE V: Object localization accuracy. The color vision effect exists in another visual recognition task like object localization, not only in classification.
 CD Group IoU-30 IoU-50 IoU-70 Easy 75.31 73.12 50.34 Medium 74.48 72.77 50.01 Hard 72.94 70.80 48.97
@@
 Results for each CD group are summarized in Tab.
-LABEL:tab:frequency.
+IV.
 We observe that when presented with only the LFC of images, models consistently perform worse in Hard group, and surprisingly, on HFC both models show better performance for Hard group.
@@
 These failure cases are deliberately selected to represent scenarios where human visual recognition remains robust, yet the models exhibit significant shortcomings.
-Accompanying each image is the corresponding attention map, we observe that the model may not be able to attend to the right object areas in whole, or sometimes the object is blended in the background.
+Accompanying each image is the corresponding attention map, though it may not be the best interpretibility technique, we observe that the model may not be able to attend to the right object areas in whole, or sometimes the object is blended in the background.
 Thus, we believe that the models’ misclassifications stem primarily from the small color difference between foreground and background elements, which appear to impede the models’ ability to effectively discern and isolate salient features.
@@
 Though the objects are easy for humans to recognize, we believe the models’ failure stem primarily from the small difference in background and foreground colors.
+IV-D Practicality of Proposed sCVR & sCR Metrics
+We want to highlight the generality of how sCVR was formulated.
+sCVR is defined to measure the variability of model performance between subgroups while taking accuracy into account.
+In this work, we have 3 different color discrimination groups per our design; however, in other scenarios, researchers may be interested in a different factor, so they have different subgroups.
+Therefore, we can design a new metric in a general way, such as sCVR to broadly assess the between-subgroup variability.
+For instance, we similarly defined an sCR (formula above) for measuring robustness to image corruption (the y-axis in Fig.
+10-left).
+We believe that solely using “total accurac” to evaluate models does not provide sufficient granularity, and a performance variability metric is needed for any study design that involves subgrouping.
+A simple statistical variance does not suffice due to subgroup sample imbalances.
+We also designed sCVR along with sCR to account for total accuracy since it is commonly regarded as the primary metric measuring model performance.
+To validate the practicality of our metric, we want to highlight that in Fig.
+10-left, our finding on model robustness to image corruption is consistent with [27] — size-wise, larger models (represented by larger ticks) have higher sCR (more robust).
+We believe that sCVR and sCR can serve as general metrics for model robustness.
+For both sCVR and sCR, we reference the definition of population standard deviation (divided by w, not w-1), rather than sample standard deviation, since our goal is to quantify a variation measurement using all the groups rather than making a specific statistical inference/estimation.
+One can also consider model sizes, but that is out of the scope of our work.
 V Other Utilities: Object Localization and Spurious Correlation
@@
 We summarize the results on ColorSenese-ImageNet in Tab.
-IV.
+V.
 We observe similar performance differences as in classification task, where the model tends to perform better as the CD group becomes easier.
@@
 The complete plot for all models is shown in Appendix.
-VI Discussion
+VI Limitations
+We study color vision in the scope of static visual perception, and it would be interesting to extend the analysis to dynamic visual perception such as video data.
+Furthermore, following our extensive analyses in Sec.
+IV-C, to address the color vision phenomenon that our work revealed in modern DL architectures, only advanced training such as data augmentation shows marginal benefit, and it would be a promising future direction to develop a bespoke data augmentation method to enhance the mitigation of color vision in modern DL architectures.
+VII Discussion
 Our work presents the development and demonstrates the utilities of our curated ColorSense dataset, which is a versatile tool for (1) investigating the impact of color vision on various visual recognition tasks, (2) featuring compatibility with OOD environmental and artificial noise datasets, and (3) controlling for spurious correlations.
@@
 Lastly, CNNs such as Convnext can be as powerful as transformers, as similarly observed in [58].
-We believe that in the era of foundation models with various scaling laws, our dataset and analysis framework are valuable and can push the boundaries of benchmarking.
-VII Potential Broader Impact
+We believe that in the era of foundation models with various scaling laws, our dataset and analysis framework are valuable and can push the boundaries of benchmarking in visual recognition.
+VIII Potential Broader Impact
 The exploration of machine-human alignment stands as a highly coveted focus in research [3].